# Degree-Weighted DeGroot Learning

### v2

@@ B.3.1 Proof of the First Part of Proposition 1
$\mathbf{T}^{\infty}_{ij}=\frac{\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{A}))}{\sum_{i,j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{A}))},\quad\forall i,j=1,\dots,n.$ (21)
$\blacksquare$
B.3.2 Proof of the Second Part of Proposition 1
The second part generalizes the result in Golub and Jackson (2012). For the proof, we establish an upper and a lower bound and show that they coincide. The main structure of the proof follows Golub and Jackson (2012), with modifications to accommodate our matrix $\mathbf{T}$ and the norm $\|\cdot\|_{2}$.
A key step in this adaptation is applying the spectral theorem to decompose $\mathbf{T}$. Lemma 1 ensures that we can apply the spectral theorem for diagonalizable matrices Meyer (2000) to $\mathbf{T}$. The spectral decomposition of $\mathbf{T}$ is given by
$\mathbf{T}=\sum_{i=1}^{n}\lambda_{i}\mathbf{U}_{i}\,,$ (22)
where $\lambda_{i}$ are the eigenvalues of $\mathbf{T}$, ordered in decreasing magnitude, and $\mathbf{U}_{i}$ are the orthogonal projections onto the corresponding eigenspaces of $\mathbf{T}$.
@@ B.3.2 Proof of the Second Part of Proposition 1
We aim to establish an upper bound for the distance $\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}$ by applying the spectral decomposition of $\mathbf{T}$ from (22). Note that
$\mathbf{T}^{t}-\mathbf{T}^{\infty}=\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\,.$ (23)
The summation starts from $i=2$ due to Lemma 2, which states that the largest eigenvalue of $\mathbf{T}$ is $1$, while all other eigenvalues have magnitudes strictly less than $1$. The term corresponding to $i=1$ cancels with $\mathbf{T}^{\infty}$ since
$\mathbf{U}_{1}=(\mathbf{v}\mathbf{w}^{\top})^{t}=\mathbf{v}(\mathbf{w}^{\top}\mathbf{v})^{t-1}\mathbf{w}^{\top}=\mathbf{v}\mathbf{w}^{\top}=\mathbf{T}^{\infty},\quad\text{for }t\in\mathbb{N}.$
Applying (23), we obtain
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}$ $=\left\lVert\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}$ (24)
$=\sum_{i=2}^{n}|\lambda_{i}|^{2t}\left\lVert\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2},\quad\quad(\mathbf{U}_{i}\mathbf{U}_{j}=0\text{ for }i\neq j)$
$\leq|\lambda_{2}|^{2t}\sum_{i=2}^{n}\left\lVert\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}$
$=|\lambda_{2}|^{2t}\left\lVert\sum_{i=2}^{n}\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}\,.\quad\quad(\mathbf{U}_{i}\mathbf{U}_{j}=0\text{ for }i\neq j)$
Define $\mathbf{U}=\sum_{i=2}^{n}\mathbf{U}_{i}$, which is a single orthogonal projection satisfying $\mathbf{U}=\mathbf{U}^{\top}=\mathbf{U}^{2}$. Then, for any vector $\mathbf{x}$,
$\|\mathbf{U}\mathbf{x}\|^{2}_{2}=\langle\mathbf{U}\mathbf{x},\mathbf{U}\mathbf{x}\rangle=\langle\mathbf{U}^{\top}\mathbf{U}\mathbf{x},\mathbf{x}\rangle=\langle\mathbf{U}\mathbf{x},\mathbf{x}\rangle\leq\|\mathbf{U}\mathbf{x}\|_{2}\|\mathbf{x}\|_{2}.$
The last inequality follows from the Cauchy-Schwarz inequality. Applying this to (24), we obtain
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}\leq|\lambda_{2}|^{2t}\|\mathbf{b}\|_{2}^{2}=|\lambda_{2}|^{2t}.$
Taking the square root on both sides yields the desired upper bound. $\blacksquare$
@@
To derive a lower bound, we consider a specific choice of $\mathbf{b}$. Let $\mathbf{b}$ be the eigenvector of $\mathbf{T}$ corresponding to its second largest eigenvalue $\lambda_{2}$ in magnitude, normalized such that $\|\mathbf{b}\|_{2}^{2}=1$. Then,
@@
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}$ $=\left\lVert\lambda_{2}^{\,t}\mathbf{b}\right\rVert^{2}_{2}$
$=|\lambda_{2}|^{2t}.$
@@
B.4 Proof of Theorem 1
B.4.1 Proof of the First Part of Theorem 1
Similar to Equation (19), we have
@@ B.4.1 Proof of the First Part of Theorem 1
Combining this with (19) and (20), we have
@@
$+\|\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{D}_{2,\mathbf{R}}^{-1/2}-\mathbf{I}\|,$ (25)
where $\mathbf{v}^{*}_{1}$ is the first eigenvector of $\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{R}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}$.
In order to control the right hand side of the inequality, we introduce the following propositions and lemmas, with their proof presented at the end of the proof of Theorem 1.
Proposition 4.
$\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\,,$
with probability at least $1-\frac{2}{n^{2}}$.
Proposition 5.
Under Assumptions (1)-(2),
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\,,$ (26)
with probability at least $1-\frac{14}{n^{2}}$.
Lemma 4.
@@
$\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{1}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (27)
and
$C_{2}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{3}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (28)
@@
$C_{1}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (29)
and
$C_{3}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{2}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (30)
Lemma 5.
For a fixed $j$,
$|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\leq\sqrt{2n\log n}\,,$ (31)
with probability at least $1-\frac{2}{n^{4}}$.
Moreover,
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|\leq\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|\,,$ (32)
for some $c_{j,1}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n},d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
We first control the second term on the right-hand side of (B.4.1). By Lemmas 45 and Proposition 5, with probability at least $1-\frac{15}{n^{2}}$, we obtain
$\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}-\mathbf{I}\|$ $\leq\|\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{D}_{\mathbf{R}}^{1/2}-\mathbf{I}\|\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}\|+\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}-\mathbf{I}\|$ (33)
@@
where $C$ is a positive constant. A similar inequality holds for the third term on the right-hand side of (B.4.1).
For the first term on the right-hand side of (B.4.1), using a similar argument as in the proof of (33) and applying the Davis-Kahan theorem along with Propositions 45 and Lemmas 45, with probability at least $1-\frac{15}{n^{2}}$, we obtain
$\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|\leq C\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}-\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\|,$ (34)
which further satisfies
$\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|\leq C\sqrt{n\log n}(\min_{i}d_{i}(\mathbf{R}))^{-1},$ (35)
where $C$ is a positive constant.
This completes the proof of the first part of Theorem 1. $\blacksquare$
B.4.2 Proof of the Second Part of Theorem 1
Recall that $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$. From the definitions of the matrices in Equations (15) and (18), and by Lemma 3, comparing the eigenvalues $\lambda_{2}(\mathbf{T})$ and $\lambda_{2}(\mathbf{T}^{*})$ reduces to comparing the eigenvalues of
$\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$
and
$\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2},$
since these matrices are similar to $\mathbf{T}$ and $\mathbf{T}^{*}$, respectively.
For simplicity, define

### v3

@@ 3.1 Proof of the First Part of Proposition 1
(B.11) $\mathbf{T}^{\infty}_{ij}=\frac{\sum_{i^{\prime}}A_{i^{\prime}j}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i^{\prime}}(\mathbf{A}))}{\sum_{i^{\prime},j^{\prime}}A_{i^{\prime}j^{\prime}}\,\phi(\alpha,d_{j^{\prime}}(\mathbf{A}))\phi(\alpha,d_{i^{\prime}}(\mathbf{A}))},\quad\forall i,j=1,\dots,n.$
$\Box$
3.2 Proof of the Second Part of Proposition 1
This part generalizes the result in Golub and Jackson (2012). For the proof, we establish an upper and a lower bound and show that they coincide. The main structure of the proof follows Golub and Jackson (2012).
A key step in this adaptation is applying the spectral theorem to decompose $\mathbf{T}$. Lemma B.1 establishes that $\mathbf{T}$ is similar to a real symmetric matrix, which allows us to apply the spectral theorem for diagonalizable matrices Meyer (2000). The spectral decomposition of $\mathbf{T}$ is given by
(B.12) $\mathbf{T}=\sum_{i=1}^{n}\lambda_{i}\mathbf{U}_{i}\,$,
where $\lambda_{i}$ are the eigenvalues of $\mathbf{T}$, ordered in decreasing magnitude, and $\mathbf{U}_{i}$ is the orthogonal projection associated with $\lambda_{i}$. Specifically, we have $\mathbf{U}_{i}=\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{v}_{i}\mathbf{v}_{i}^{T}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}$, where $\mathbf{v}_{i}$ is the unit eigenvector of the symmetric matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$ associated with $\lambda_{i}$.
@@ 3.2 Proof of the Second Part of Proposition 1
We aim to establish an upper bound for the distance $\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{\mathbf{s}}$ by applying the spectral decomposition of $\mathbf{T}$ from (B.12). Note that
(B.13) $\mathbf{T}^{t}-\mathbf{T}^{\infty}=\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\,.$
The summation starts from $i=2$ due to Lemma B.2, which states that the largest eigenvalue of $\mathbf{T}$ is $1$, while all other eigenvalues have magnitudes strictly less than $1$. The term corresponding to $i=1$ cancels with $\mathbf{T}^{\infty}$ since $\mathbf{U}_{1}=(\mathbf{v}\mathbf{w}^{\top})^{t}=\mathbf{v}(\mathbf{w}^{\top}\mathbf{v})^{t-1}\mathbf{w}^{\top}=\mathbf{v}\mathbf{w}^{\top}=\mathbf{T}^{\infty},\quad\text{for }t\in\mathbb{N}.$ Applying (B.13), we obtain
(B.14) $\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{\mathbf{s}}^{2}$ $=\left\lVert\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{\mathbf{s}}$
$=\sum_{i=2}^{n}|\lambda_{i}|^{2t}\left\lVert\mathbf{v}_{i}\mathbf{v}_{i}^{T}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{b}\right\rVert^{2}_{2}/D,\quad\quad(\mathbf{v}_{i}\mathbf{v}_{j}=0\text{ for }i\neq j)$
$\leq|\lambda_{2}|^{2t}\sum_{i=2}^{n}\left\lVert\mathbf{v}_{i}\mathbf{v}_{i}^{T}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{b}\right\rVert^{2}_{2}/D$
$=|\lambda_{2}|^{2t}\left\lVert\sum_{i=2}^{n}\mathbf{v}_{i}\mathbf{v}_{i}^{T}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{b}\right\rVert^{2}_{2}/D\,.\quad\quad(\mathbf{v}_{i}\mathbf{v}_{j}=0\text{ for }i\neq j)$
where $D=\sum\limits_{i^{\prime},j}A_{i^{\prime}j}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i^{\prime}}^{\alpha}(\mathbf{A})$. Noticing that $\sum_{i=2}^{n}\mathbf{v}_{i}\mathbf{v}_{i}^{T}\leq\mathbf{I}$, we obtain
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{\mathbf{s}}^{2}\leq|\lambda_{2}|^{2t}\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{b}\|_{2}^{2}/D=|\lambda_{2}|^{2t}|\|\mathbf{b}\|_{\mathbf{s}}^{2}=|\lambda_{2}|^{2t}.$
Taking the square root on both sides yields the desired upper bound. $\Box$
@@
To derive a lower bound, we consider a specific choice of $\mathbf{b}$—the eigenvector of $\mathbf{T}$ corresponding to $\lambda_{2}$ in magnitude, normalized such that $\|\mathbf{b}\|_{\mathbf{s}}^{2}=1$. Then
@@
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{\mathbf{s}}^{2}$ $=\left\lVert\lambda_{2}^{\,t}\mathbf{b}\right\rVert^{2}_{\mathbf{s}}=|\lambda_{2}|^{2t}.$
@@
4 Proof of Theorem 1
4.1 Proof of the First Part of Theorem 1
Similar to Equation (B.9), we have
@@ 4.1 Proof of the First Part of Theorem 1
Combining this with (B.9) and (B.10), we have
@@
(B.15) $+\|\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{D}_{2,\mathbf{R}}^{-1/2}-\mathbf{I}\|,$
where $\mathbf{v}^{*}_{1}$ is the first eigenvector of $\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}$.
In order to control the right-hand side of the inequality, we introduce the following propositions and lemmas, with their proofs presented in Online Appendix O1
Proposition B.1
$\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\,$, with probability at least $1-{2}/{n^{2}}$.
Proposition B.2
Under Assumptions 1-2, with probability at least $1-{14}/{n^{2}}$,
(B.16) $\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\,.$
Lemma B.4
@@
(B.17) $\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{1}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,$,
(B.18) $C_{2}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{3}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,$,
@@
(B.19) ${C_{4}}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,$,
(B.20) ${C_{5}}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq{C_{6}}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,$,
where $C_{4},C_{5},C_{6}$ are positive constants independent of $n$.
Lemma B.5
For a fixed $j$, with probability at least $1-{2}/{n^{4}}$,
(B.21) $|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\leq\sqrt{2n\log n}\,.$
Moreover, for some $c_{j,1}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n},d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$, j=1,,nj=1,...,n, with probability at least $1-{2}/{n^{4}}$,
(B.22) $\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|\leq\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|\,.$
We first control the second term on the right-hand side of (4.1). By Lemmas B.4B.5 and Proposition B.2, with probability at least $1-{15}/{n^{2}}$, we obtain, for some positive constant $C$,
(B.23) $\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}-\mathbf{I}\|$ $\leq\|\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{D}_{\mathbf{R}}^{1/2}-\mathbf{I}\|\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}\|+\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}-\mathbf{I}\|$
@@
A similar inequality holds for the third term on the right-hand side of (4.1).
For the first term on the right-hand side of (4.1), applying the Davis-Kahan theorem (Davis and Kahan, 1970; Yu, Wang and Samworth, 2015) and using Assumption 3 (which ensures that the spectral gap $1-\lambda_{2}(\mathbf{T}^{*})$ is eventually bounded below by some positive constant 1c>01-c>0), we obtain
(B.24) $\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|\leq C\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}-\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\|,$
where $C$ is a positive constant inversely related to the spectral gap 1c>01-c>0. Furthermore, using a similar argument as in the proof of (B.23), and applying Propositions B.1B.2 and Lemmas B.4B.5 to control the differences $\|\mathbf{D}_{2,\mathbf{A}}^{1/2}-\mathbf{D}_{2,\mathbf{R}}^{1/2}\|$, $\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}-\mathbf{D}_{1,\mathbf{R}}^{-1/2}\|$, and $\|\mathbf{A}-\mathbf{R}\|$, we find that with probability at least $1-{15}/{n^{2}}$,
(B.25) $\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}-\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\|\leq\tilde{C}\sqrt{n\log n}(\min_{i}d_{i}(\mathbf{R}))^{-1}=\tilde{C}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}$
for some $\tilde{C}>0$. This completes the proof of the first part of Theorem 1. $\Box$
4.2 Proof of the Second Part of Theorem 1
Recall that $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$. From the definitions of the matrices in Equations (B.5) and (B.8), and by Lemma B.3, comparing the eigenvalues $\lambda_{2}(\mathbf{T})$ and $\lambda_{2}(\mathbf{T}^{*})$ reduces to comparing the eigenvalues of $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$ and $\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2},$ since these matrices are similar to $\mathbf{T}$ and $\mathbf{T}^{*}$, respectively. For simplicity, define